# MPCM-Net: Multi-scale network integrates partial attention convolution with Mamba for ground-based cloud image segmentation

### v1

@@ III-C2 Partial Spatial Module
While downsampling operations across successive encoder layers inevitably degrade spatial location information in feature maps, this property paradoxically mitigates boundary ambiguity in ground-based cloud segmentation tasks. As illustrated in Fig. 3 (c), we adapt the ParCM paradigm by integrating partial channel information with an attention mechanism and a 1 $\times$ 1 convolution for effective channel mixing. Unlike conventional spatial attention, relying solely on pointwise convolution, our approach employs a spatial selectivity mechanism to capture precise positional information. This mechanism utilizes spatial feature descriptors to focus on the most relevant spatial regions adaptively. Specifically, given input feature map $C_{2}\in\mathbb{R}^{C\times\left(1-R\right)\times H\times W}$, we apply global max pooling and global average pooling to derive two complementary feature maps enriched with detailed spatial information. These maps are concatenated along the channel dimension to enhance global context. A convolutional operation reduces the fused feature dimensionality. A sigmoid activation produces the spatial attention descriptor $\tilde{C}$ for weighting multi-scale features to obtain the final output $\hat{C}$. The specific formula for this process is as follows:
$\tilde{C}$ $=\sigma\left(\mathrm{conv}\left(\left[\mathrm{GMP}\left(C_{2}\right),\mathrm{GAP}\left(C_{2}\right)\right]\right)\right)$ (7)
$\hat{C}$ $=\tilde{C}\ast C_{2}$ (8)
where $\mathrm{GAP}\left(\cdot\right)$ and $\mathrm{GMP}\left(\cdot\right)$ are the global max pooling and global average pooling, respectively, and $\left[\cdot,\cdot\right]$ denotes the concatenation of the feature maps. The output of the ParSM is $S_{x}$:
[…]
@@ III-C3 Partial Attention Module
Employing self-attention in the final layer of MPAC provides an effective method to capture discriminative features and global dependencies. This approach not only facilitates spatial information interaction but also expands the effective receptive field of the model. An appropriately designed attention mechanism significantly enhances segmentation precision for multi-scale cloud clusters in ground-based imagery.
Vision Transformer (ViT), a prominent research focus in computer vision, utilizes softmax attention as its core operator. Although ViTs demonstrate superior performance in semantic segmentation tasks, including ground-based cloud segmentation, they suffer from quadratic computational complexity $O$ ($N^{2}$) relative to input sequence length $N$, resulting in prohibitive computational costs. Linear attention mechanism addresses this limitation by reformulating the computation order of Query-Key-Value ($Q$, $K$, and $V$) operations and eliminating softmax, thereby reducing complexity to $O$ (NN). However, this efficiency gain comes at the expense of substantial performance degradation.
Inspired by MALA, we introduce an amplitude-aware mechanism into the computation logic of the linear attention, proposing a novel Softmax-like Linear Attention (SLA) mechanism. SLA preserves the amplitude propagation capability characteristic of softmax attention while maintaining linear complexity, thereby emulating the dynamic distribution properties of conventional attention scores. When integrated into the ParAM as a replacement for standard self-attention, SLA effectively combines the high expressiveness of softmax attention with the computational efficiency of linear attention, achieving dual advantages in performance and complexity.
The performance degradation in linear attention mechanisms stems from their inability to propagate Q-amplitude values. Consequently, attention scores across different features remain confined to an identical distribution, failing to dynamically focus on salient regions in response to variations in Q-values. The core innovation of SLA addresses this limitation by introducing an offset term to amplify weights associated with higher original attention scores, thereby enhancing local feature concentration. Specifically, as illustrated in Fig. 3 (g), we reformulate the similarity computation by incorporating a scaling factor $\alpha$ and offset $\beta$ into the original linear attention framework, formulating a novel attention metric as follows:
@@
$\begin{split}\mathrm{Att}_{i}&=\sum_{j=1}^{N}\frac{\phi\left(Q_{i}\right)\phi\left(K_{j}\right)^{T}}{\sum_{j=1}^{N}\phi\left(Q_{i}\right)\phi\left(K_{j}\right)^{T}}V_{j}\\ &=\sum_{j=1}^{N}\frac{li}{\sum_{j=1}^{N}li}V_{j}\end{split}$ (11)
$\begin{split}\mathrm{SLA}_{i}&=\sum_{j=1}^{N}\left[\frac{li}{\sum_{j=1}^{N}li}\left(1+\sum_{j=1}^{N}li\right)-\frac{\sum_{j=1}^{N}li}{N}\right]V_{j}\\ &=\sum_{j=1}^{N}\left[\left(1+\frac{1}{\sum_{j=1}^{N}li}\right)li-\frac{\sum_{j=1}^{N}li}{N}\right]V_{j}\\ &=\sum_{j=1}^{N}\left[\alpha\cdot li-\beta\right]V_{j}\end{split}$ (12)
where $\mathrm{Att}_{i}$ denotes the original linear attention with an input of $N$ tokens which represented as $x\in\mathbb{R}^{N\times C}$, $li$ denotes $\phi\left(Q_{i}\right)\phi\left(K_{j}\right)^{T}$ and $\mathrm{SLA}_{i}$ denotes the SLA. $\alpha$ denotes the scale factor and $\beta$ denotes the offset term.
Given $Q_{i}$, the ratio of its attention scores toward two distinct keys $K_{m}$ and $K_{n}$ can be expressed as follows:
$\frac{\alpha\operatorname{Sim}\left(Q_{i},K_{m}\right)-\beta}{\alpha\operatorname{Sim}\left(Q_{i},K_{n}\right)-\beta}=w$ (13)
Assuming $Q$ assigns a higher attention score to $K_{m}$, when the magnitude of $Q$ scales by a factor $s>1$, the updated values of alphaalpha and betabeta can be derived as:
$\begin{split}\alpha_{up}&=1+\frac{1}{s\sum_{j=1}^{N}ex}\\ &=1+\frac{1}{s}\left(\alpha-1\right)=\frac{s+\alpha-1}{s}\end{split}$ (14)
$\begin{split}\beta_{up}=\frac{s\sum_{j=1}^{N}ex}{N}=s\beta\end{split}$ (15)
@@
$\begin{split}w_{up}&=\frac{\alpha_{up}\operatorname{Sim}\left(Q_{i},K_{m}\right)-\beta_{up}}{\alpha_{up}\operatorname{Sim}\left(Q_{i},K_{n}\right)-\beta_{up}}\\ &=\frac{\left(s+\alpha-1\right)\operatorname{Sim}\left(Q_{i},K_{m}\right)-s\beta}{\left(s+\alpha-1\right)\operatorname{Sim}\left(Q_{i},K_{n}\right)-s\beta}\\ &=\frac{\alpha\operatorname{Sim}\left(Q_{i},K_{m}\right)-\frac{s\alpha}{s+\alpha-1}\beta}{\alpha\operatorname{Sim}\left(Q_{i},K_{n}\right)-\frac{s\alpha}{s+\alpha-1}\beta}\end{split}$ (16)
Since $s>1$, $a>1$, and $b>1$, it can be proven that $w_{up}>w$. This demonstrates that SLA allocates greater attention to features originally receiving higher scores while reducing attention to features with lower scores, thereby successfully emulating the strong representational capacity of standard softmax attention.
@@ III-D Multi Mamba Decoding
Establishing global dependencies over long-range features enables the extraction of comprehensive cloud cluster information, thus mitigating the loss of boundary details in ground-based cloud imagery. Existing decoders primarily utilize lateral connections to fuse same-scale features from the encoder, compensating for local information loss during downsampling, while employing various attention mechanisms to enhance feature representation. However, this approach suffers from several limitations: 1) Transformer-based decoders significantly increase computational complexity, adversely affecting the inference efficiency of ground-based cloud segmentation; 2) The upsampling process lacks global feature correlations between local information across hierarchical levels. In fact, Mamba [Mamba] has been shown to establish global context while maintaining linear time complexity. Nevertheless, its compressed fixed-dimensional state representation can lead to partial loss of local information.
To this end, we propose a decoder architecture incorporating a Multi-scale Mamba Block. This design enhances the capacity of the model for capturing global dependencies among local features across hierarchical levels, compensates for lost contextual information between the encoder and decoder, and improves boundary segmentation in ground-based cloud imagery. Specifically, we design a multi-scale feature enhancement module based on the Mamba architecture within the decoder by introducing a spatial-semantic hybrid domain to process feature maps at varying granularities. The module achieves dynamic and adaptive extraction of multi-scale semantic information. Collectively, our decoder performs deep feature aggregation both spatially and across scales, significantly mitigating information loss caused by translucent cloud boundaries in ground-based cloud images.
@@
Figure 5: (a) The structure of the proposed SSHD is composed of two parts: $X_{C\times\left(1-R\right)}$ and $X_{C\times R}$. The $X_{C\times\left(1-R\right)}$ passes through the 2D-SSM, the $X_{C\times R}$ passes through the HA module, and obtains the SSHD output $X_{s}$. The bottom right side of the figure shows the notes of some modules.
@@ III-D1 Multi-scale Mamba Block
To mitigate boundary information loss in ground-based cloud imagery and establish long-range global dependencies, we propose a Mamba-based module. As illustrated in Fig. 4, to achieve multi-scale processing within the Mamba framework, we first employ large-kernel receptive field aggregation on $f_{d}$ to extract features at varying scales. Specifically, feature map $f_{d}$ is processed with a 3 $\times$ 3 convolution (stride = 2), aggregating neighborhood features within a 9 - pixel receptive field, yielding $f_{d5}$ with dimensions $H/16\times W/16\times 2C$. Feature map fd is processed with a 5 $\times$ 5 convolution (stride = 4), aggregating neighborhood features within a 25 - pixel receptive field, yielding $f_{d4}$ with dimensions $H/32\times W/32\times 4C$. The multi-scale feature maps are subjected to a pixel unshuffle procedure in order to allow the Mamba structure to progressively scan the channel dimensions of $f_{d}$, $f_{d3}$, and $f_{d5}$. This ensures uniform spatial dimensions across all three features while minimizing information loss inherent in downsampling. The core idea involves spatially rearranging $f_{d}$ and $f_{d3}$: non-overlapping spatial blocks are shifted into the channel dimension, preserving the spatial integrity of the original resolution while matching the reduced dimensions of $f_{d5}$, resulting in transformed features $f_{d1}^{{}^{\prime}}$ and $f_{d3}^{{}^{\prime}}$. The following equations formally describe this transformation:
@@
where $\mathrm{PU}_{i\times j}\left(\cdot\right)$ denotes a pixel unshuffle operation using non-overlapping $i\times j$ image patches. Subsequently, the three feature maps are projected to a uniform channel dimension via 1 $\times$ 1 convolutions and concatenated along the channel axis. This concatenated feature is then fed into the proposed Spatial-Semantic Hybrid Domain module to enhance local spatial features and establish global semantic dependencies. The resulting feature map possesses $H/32\times W/32\times 3C$, where the three segments along the channel dimension correspond to the input features $F_{d1}^{{}^{\prime}}$, $F_{d3}^{{}^{\prime}}$, and $F_{d4}^{{}^{\prime}}$, respectively. To maintain consistent spatial resolution throughout the decoder, the hybrid features output by SSHD are upsampled to match the spatial scale of $u_{2}$ via bilinear interpolation. Finally, a 1 $\times$ 1 convolution reduces dimensionality, yielding the multi-scale global feature $f_{s}$, which integrates local feature relationships.
@@ III-D2 Spatial-Semantic Hybrid Domain
Decoders widely adopted in prior ground-based cloud segmentation methods typically employ self-attention mechanisms to capture long-range dependencies within cloud clusters. However, this approach inherently suffers from quadratic computational complexity, presenting significant challenges for real-time inference requirements in ground-based cloud analysis. Motivated by the success of Mamba in maintaining linear complexity while modeling global context, we propose a SSHD module based on the Mamba architecture to enhance global feature representation of cloud formations.
The proposed SSHD module, illustrated in Fig. 5 (a), processes input features similarly to the MPAC within the decoder. First, input feature $X\in\mathbb{R}^{C\times H\times R}$ undergoes channel partitioning with ratio $R$ (hyperparameter $R$ = 1/4, consistent with MPAC). The partitioned component $X_{C\times\left(1-R\right)}$ is then processed sequentially through a linear layer and a depth-wise convolutional (DW-Conv) layer to enhance feature representation. Subsequently, 2D State Space Model (2D-SSM) [Vmamba] extraction captures global contextual information incorporating local features, followed by layer normalization to genearte $X_{1}$. To extract fine-grained information from multi-level features, a parallel branch $X_{C\times R}$ introduces Hybrid Attention (HA), integrating channel and spatial attention mechanisms, for feature refinement to generate $X_{2}$. Finally, features from both branches are concatenated along the channel dimension and processed through a linear layer, yielding the SSHD output $X_{s}$. The complete workflow is formally described by the following equations:
$X_{1}$ $=\mathrm{LN}\left(\mathrm{SSM}\left(\mathrm{DW}\left(\mathrm{Linear}\left(X_{C\times\left(1-R\right)}\right)\right)\right)\right)$ (19)
$X_{2}$ $=\mathrm{HA}\left(X_{C\times R}\right)$ (20)
$X_{s}$ $=\mathrm{Linear}\left(X_{1}\oplus X_{2}\right)$ (21)
where $\mathrm{DW}\left(\cdot\right)$ denotes depth-wise convolution, $\mathrm{HA}\left(\cdot\right)$ denotes the hybrid attention, $\mathrm{SSM}\left(\cdot\right)$ denotes the 2D selective scan module, and $\oplus$ denotes element-wise production.
While the Mamba model was originally designed for natural language processing tasks using sequential selective scanning of 1D data, its direct application to 2D image processing presents inherent challenges. Inspired by [Vmamba], we introduce a 2D selective scanning module for ground-based cloud image segmentation. As illustrated in Fig. 5 (b), the 2D feature map $X_{C\times\left(1-R\right)}$ is first flattened into a 1D sequence, then selectively scanned along four spatial directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. During each directional scan, the selective state space model efficiently captures orientation-specific global contextual information and establishes long-range dependencies. Finally, the scanned sequences from all directions are merged to reconstruct the original 2D feature representation.
@@
We propose HA to enhance the interaction between spatial and semantic information in the MMD, strengthening the capability of the model for extracting cloud boundary features in ground-based cloud imagery, as illustrated in Fig.  6. Specifically, the feature subset $X_{C\times R}$ first undergoes channel attention for spatial information interaction. Since feature maps exhibit approximately Gaussian distributions during training, whereas standard Squeeze-and-Excitation (SE) blocks utilize only mean statistics, we introduce an enhanced Gaussian-SE module that compresses global spatial information using both mean and variance to model Gaussian statistics for channel-wise recalibration. Subsequently, spatial attention is implemented via point-wise convolution, compressing global channel information into a single-channel tensor. These tensors are processed through a Hard-Sigmoid activation function to generate the spatial attention map for feature weighting. The complete workflow is formally described by the following equations:
$\mathrm{HA}$ $=\mathrm{S_{A}}\left(\mathrm{G_{SE}}\left(X_{C\times R}\right)\right)$ (22)
$\mathrm{G_{SE}}$ $=\sigma_{H}\left(\mathrm{LR}\left[\mathrm{GAP}\left(X_{C\times R}\right),\mathrm{STD}\left(X_{C\times R}\right)\right]\right)$ (23)
@@
where $\mathrm{S_{A}}\left(\cdot\right)$ denotes spatial attention, $\mathrm{G_{SE}}$ denotes Gaussian-SE module, $\mathrm{GAP}\left(\cdot\right)$, $\mathrm{STD}\left(\cdot\right)$, and $\mathrm{GMP}\left(\cdot\right)$ are the global max pooling, standard deviation, and global average pooling, respectively, $\sigma_{H}\left(\cdot\right)$ is the Hard-Sigmoid activation function, and $\left[\cdot,\cdot\right]$ denotes the concatenation of the feature maps.
Finally, $f_{s}$ is upsampled and concatenated with the encoder’s first-layer feature. The resulting feature map undergoes a standard convolution to restore the input resolution, followed by a 1 $\times$ 1 convolution for channel adjustment prior to final prediction.

### v2

@@ III-C2 Partial Spatial Module
While downsampling operations across successive encoder layers inevitably degrade spatial location information in feature maps, this property paradoxically mitigates boundary ambiguity. As illustrated in Fig. 3 (c), firstly, the input features are divided into two sub-features, $I_{prim}\in\mathbb{R}^{C_{prim}\times H\times W}$ and $I_{sec}\in\mathbb{R}^{C_{sec}\times H\times W}$, by a ratio $r$. Then we adapt the ParCM paradigm by integrating partial channel information with an attention mechanism and a 1 $\times$ 1 convolution for effective channel mixing. Unlike conventional spatial attention, relying solely on pointwise convolution, our approach employs a spatial selectivity mechanism to capture precise positional information. This mechanism utilizes spatial feature descriptors to focus on the most relevant spatial regions adaptively. Specifically, given input feature map $I_{sec}$, we apply global max pooling and global average pooling to derive two complementary feature maps enriched with detailed spatial information. These maps are concatenated along the channel dimension to enhance global context. A convolutional operation reduces the fused feature dimensionality. A sigmoid activation produces the spatial attention descriptor $\tilde{I}$ for weighting multi-scale features to obtain the final output $\hat{I}$. The specific formula for this process is as follows:
$\tilde{I}$ $=\sigma\left(\mathrm{conv}\left(\left[\mathrm{GMP}\left(I_{sec}\right),\mathrm{GAP}\left(I_{sec}\right)\right]\right)\right)$ (6)
$\hat{I}$ $=\tilde{I}\ast I_{sec}$ (7)
where $\mathrm{GAP}\left(\cdot\right)$ and $\mathrm{GMP}\left(\cdot\right)$ are the global average pooling and global max pooling, respectively. The output of the ParSM is $S_{x}$:
$S_{x}=\left[\mathrm{conv}\left(I_{prim}\right),\hat{I}\right$] (8)
@@ III-C3 Partial Attention Module
In the final layer of MPAC, to adaptively capture global semantic dependencies, we propose ParAM that adopts a “split-and-process” strategy consistent with the design philosophy of ParCM and ParSM.
As illustrated in Fig. 3 (f), we first split the input feature map. The sub-feature $X_{sec}$ is processed by our proposed Softmax-like Linear Attention (SLA) (The subsequent text will provide a comprehensive description) to focus on extracting long-range contextual dependencies. Finally, the outputs from two branches are concatenated.
An appropriately designed attention mechanism in ParAM significantly enhances segmentation precision for multi-scale cloud clusters. ViTs suffer from quadratic computational complexity $O$ ($N^{2}$) relative to input sequence length NN[10]. Linear attention addresses this limitation by reformulating the computation order of Query-Key-Value operations and eliminating softmax, thereby reducing complexity to OO(NN) [25]. However, this efficiency gain comes at the expense of substantial performance degradation. Inspired by MALA [12], we introduce an amplitude-aware mechanism into the computation logic of the linear attention, proposing a novel Softmax-like Linear Attention mechanism. SLA preserves the amplitude propagation capability characteristic of softmax attention.
The performance degradation in linear attention mechanisms stems from their inability to propagate Q-amplitude values. Consequently, attention scores across different features remain confined to an identical distribution, failing to dynamically focus on salient regions in response to variations in Q-values. This mechanism utilizes the kernel function $\phi$ (.) to approximate the similarity function and maps $Q$ and $K$ to positive real numbers. The core innovation of SLA addresses this limitation by introducing an offset term to amplify weights associated with higher original attention scores, thereby enhancing local feature concentration. Specifically, as illustrated in Fig. 3 (g), we reformulate the similarity computation by incorporating a scaling factor $\alpha$ and offset $\beta$ into the original linear attention framework, formulating a novel attention metric as follows:
@@
$\begin{split}\mathrm{Att}_{i}&=\sum_{j=1}^{N}\frac{\phi\left(Q_{i}\right)\phi\left(K_{j}\right)^{T}}{\sum_{m=1}^{N}\phi\left(Q_{i}\right)\phi\left(K_{m}\right)^{T}}V_{j}\\ &=\sum_{j=1}^{N}\frac{l_{ij}}{\sum_{m=1}^{N}l_{im}}V_{j}\end{split}$ (10)
$\begin{split}\mathrm{SLA}_{i}&=\sum_{j=1}^{N}\left[\frac{l_{ij}}{\sum_{m=1}^{N}l_{im}}\left(1+\sum_{m=1}^{N}l_{im}\right)-\frac{\sum_{m=1}^{N}l_{im}}{N}\right]V_{j}\\ &=\sum_{j=1}^{N}\left[\left(1+\frac{1}{\sum_{m=1}^{N}l_{im}}\right)l_{ij}-\frac{\sum_{m=1}^{N}l_{im}}{N}\right]V_{j}\\ &=\sum_{j=1}^{N}\left(\alpha\cdot l_{ij}-\beta\right)V_{j}\end{split}$ (11)
where $\mathrm{Att}_{i}$ denotes the original linear attention with an input of $N$ tokens which represented as $x\in\mathbb{R}^{N\times C}$. $l_{\left(xy\right)}$ denotes $\phi\left(Q_{x}\right)\phi\left(K_{y}\right)^{T}$. $\phi(\cdot)$ denotes the kernel feature map function. We use $\phi(x)=\mathrm{ELU}\left(x\right)+1$, where $\mathrm{ELU}$ denotes the Exponential Linear Unit: $\mathrm{ELU}\left(x\right)=e^{x}-1$ for $\leq 0, xx$ for $x>0$. The +1+1 shift ensures that all kernel values are strictly positive, which is a mathematical prerequisite for the validity of the subsequent attention probability mass function [25]. $\alpha=\left(1+1/\sum_{m=1}^{N}l_{im}\right)$ denotes the scale factor and $\beta=1/N\cdot\sum_{m=1}^{N}l_{im}$ denotes the offset term.
Given $Q_{i}$, the ratio of its attention scores toward two distinct keys $K_{n}$ and $K_{t}$ can be expressed as follows:
$\frac{\alpha l_{in}-\beta}{\alpha l_{it}-\beta}=w$ (12)
Assuming $Q$ assigns a higher attention score to $K_{n}$, when the magnitude of $Q$ scales by a factor $s>1$, the updated values of $\alpha$ and $\beta$ can be derived as:
$\begin{split}\alpha_{up}&=1+\frac{1}{s\sum_{m=1}^{N}l_{im}}\\ &=1+\frac{1}{s}\left(\alpha-1\right)=\frac{s+\alpha-1}{s}\end{split}$ (13)
$\begin{split}\beta_{up}=\frac{s\sum_{m=1}^{N}l_{im}}{N}=s\beta\end{split}$ (14)
@@
$\begin{split}w_{up}&=\frac{\alpha_{up}l_{in}-\beta_{up}}{\alpha_{up}l_{it}-\beta_{up}}\\ &=\frac{\left(s+\alpha-1\right)l_{in}-s\beta}{\left(s+\alpha-1\right)l_{it}-s\beta}\\ &=\frac{\alpha l_{in}-\frac{s\alpha}{s+\alpha-1}\beta}{\alpha l_{it}-\frac{s\alpha}{s+\alpha-1}\beta}\end{split}$ (15)
Since $s>1$, $\alpha>1$, and $\frac{s\alpha}{s+\alpha-1}>1$, it can be proven that $w_{up}>w$.
Proof of $w_{up}>w$.
We define $A=\alpha l_{in}$, $B=\alpha l_{it}$ and $k=\frac{s\alpha}{s+\alpha-1}$, where $s\alpha-(s+\alpha-1)=(s-1)(\alpha-1)>0$. We obtain $s\alpha>s+\alpha-1$, so $k>1$. We assume that $Q_{i}$ assigns a higher score to $K_{n}$, thus $A>B$. Consider the function $f(x)=\frac{A-x}{B-x}$. Its derivative with respect to $x$ is: $f^{\prime}(x)=\frac{A-B}{(B-x)^{2}}$. Since $A>B$, $f^{\prime}(x)>0$, implying $f(x)$ is strictly increasing. Since $k\beta>\beta$, it follows that $f(k\beta)>f(\beta)$, which proves $w_{up}>w$.
This demonstrates that SLA allocates greater attention to features originally receiving higher scores while reducing attention to features with lower scores, thereby successfully emulating the strong representational capacity of standard softmax attention.
Figure 5: (a) The structure of the proposed SSHD is composed of two parts: $X_{sec}$ and $X_{prim}$. The $X_{sec}$ passes through the 2D-SSM, the $X_{prim}$ passes through the HA module, and obtains the SSHD output $X_{s}$. (b) The specific structure of 2D-SSM. The 2D feature map is first flattened into a 1D sequence, then the orientation-specific information is captured via a selective state space model. Finally, the scanned sequences are merged to reconstruct the original 2D feature. The right side of the figure shows the notes of some modules.
@@ III-D Multi Mamba Decoding
Establishing global dependencies over long-range features enables the extraction of comprehensive cloud cluster information. Existing computationally expensive decoders lack global feature correlations between local information across hierarchical levels. In fact, Mamba [15] has been shown to establish global context. Nevertheless, its compressed fixed-dimensional state representation can lead to partial loss of local information.
To this end, we propose a decoder architecture incorporating a Multi-scale Mamba Block. Specifically, we design a multi-scale feature enhancement module based on the Mamba architecture within the decoder by introducing a spatial-semantic hybrid domain to process feature maps at varying granularities. The module achieves dynamic and adaptive extraction of multi-scale semantic information. Collectively, our decoder performs deep feature aggregation both spatially and across scales, significantly mitigating information loss caused by translucent cloud boundaries.
@@ III-D1 Multi-scale Mamba Block
As illustrated in Fig. 4, to achieve multi-scale processing within the Mamba framework, we first employ large-kernel receptive field aggregation on $f_{d}$ to extract features at varying scales. Specifically, feature map $f_{d}$ is processed with a 3 $\times$ 3 convolution (stride = 2), aggregating neighborhood features within a 9 - pixel receptive field, yielding $f_{d3}$ with dimensions $H/16\times W/16\times 2C$. Feature map $f_{d}$ is processed with a 5 $\times$ 5 convolution (stride = 4), aggregating neighborhood features within a 25 - pixel receptive field, yielding $f_{d5}$ with dimensions $H/32\times W/32\times 4C$. The multi-scale feature maps are subjected to a pixel unshuffle procedure in order to allow the Mamba structure to progressively scan the channel dimensions of $f_{d}$, $f_{d3}$, and $f_{d5}$. This ensures uniform spatial dimensions across all three features while minimizing information loss inherent in downsampling. The core idea involves spatially rearranging $f_{d}$ and $f_{d3}$: non-overlapping spatial blocks are shifted into the channel dimension, preserving the spatial integrity of the original resolution while matching the reduced dimensions of $f_{d5}$, resulting in transformed features $f_{d1}^{{}^{\prime}}$ and $f_{d3}^{{}^{\prime}}$. The following equations formally describe this transformation:
@@
where $\mathrm{PU}_{i\times j}\left(\cdot\right)$ denotes a pixel unshuffle operation using non-overlapping $i\times j$ image patches. Subsequently, the three feature maps are projected to a uniform channel dimension via 1 $\times$ 1 convolutions and concatenated along the channel axis. This concatenated feature is then fed into the proposed SSHD to enhance local and global features. The resulting feature map possesses $H/32\times W/32\times 3C$, where the three segments along the channel dimension correspond to the input features $F_{d1}^{{}^{\prime}}$, $F_{d3}^{{}^{\prime}}$, and $F_{d4}^{{}^{\prime}}$, respectively. To maintain consistent spatial resolution throughout the decoder, the hybrid features output by SSHD are upsampled to match the spatial scale of $u_{2}$ via bilinear interpolation. Finally, a 1 $\times$ 1 convolution reduces dimensionality, yielding the multi-scale global feature $f_{s}$, which integrates local feature relationships.
@@ III-D2 Spatial-Semantic Hybrid Domain
Previous methods present significant challenges for real-time inference requirements in ground-based cloud analysis. Motivated by the success of Mamba, we propose a SSHD module based on the Mamba architecture to enhance global feature representation.
The proposed SSHD module, illustrated in Fig. 5 (a), processes input features similarly to the MPAC within the decoder. First, input feature $X\in\mathbb{R}^{C\times H\times W}$ undergoes channel partitioning with ratio $r$. The partitioned component $X_{sec}$ is then processed sequentially through a linear layer and a depth-wise convolutional (DW-Conv) layer to enhance feature representation. Subsequently, 2D State Space Model (2D-SSM) [31] captures global contextual information incorporating local features, followed by layer normalization to genearte $X_{1}$. To extract fine-grained information, a parallel branch $X_{prim}$ introduces Hybrid Attention (HA), integrating channel and spatial attention mechanisms, for feature refinement to generate $X_{2}$. Finally, features from both branches are concatenated along the channel dimension and processed through a linear layer, yielding the SSHD output $X_{s}$. The complete workflow is formally described by the following equations:
$X_{1}$ $=\mathrm{LN}\left(\mathrm{SSM}\left(\mathrm{DW}\left(\mathrm{Linear}\left(X_{sec}\right)\right)\right)\right)$ (18)
$X_{2}$ $=\mathrm{HA}\left(X_{prim}\right)$ (19)
$X_{s}$ $=\mathrm{Linear}\left(\left[X_{1},X_{2}\right]\right)$ (20)
where $\mathrm{LN}(\cdot)$ denotes Layer Normalization, and $\mathrm{Linear}(\cdot)$ denotes a fully connected projection layer.
Inspired by [31], we introduce a 2D selective scanning module. As illustrated in Fig. 5 (b), the 2D feature map $X_{sec}$ is first flattened into a 1D sequence, then selectively scanned along four spatial directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. During each directional scan, the selective state space model efficiently captures orientation-specific global contextual information and establishes long-range dependencies. Finally, the scanned sequences from all directions are merged to reconstruct the original 2D feature representation.
@@
We propose HA to enhance the interaction between spatial and semantic information in the MMD, as illustrated in Fig.  6. Specifically, the feature subset $X_{prim}$ first undergoes channel attention. Since feature maps exhibit approximately Gaussian distributions during training, whereas standard Squeeze-and-Excitation blocks utilize only mean statistics [21], we introduce an enhanced Gaussian-SE module that compresses global spatial information using both mean and variance to model Gaussian statistics for channel-wise recalibration. Subsequently, spatial attention is implemented via point-wise convolution, compressing global channel information into a single-channel tensor. These tensors are processed through a Hard-Sigmoid activation function to generate the spatial attention map for feature weighting. The complete workflow is formally described by the following equations:
$\mathrm{HA}$ $=\mathrm{S_{A}}\left(\mathrm{G_{SE}}\left(X_{prim}\right)\right)$ (21)
$\mathrm{G_{SE}}$ $=\sigma_{H}\left(\mathrm{ReLU}(\mathrm{Linear}(\mathrm{GAP}\left(X_{prim}\right),\mathrm{STD}\left(X_{prim}\right)))\right)$ (22)
@@
where $\mathrm{S_{A}}\left(\cdot\right)$ denotes spatial attention, $\mathrm{G_{SE}}$ denotes Gaussian-SE module, $\mathrm{STD}\left(\cdot\right)$ denotes the standard deviation, $X_{in}$ denotes the input of SA, and $\sigma_{H}\left(\cdot\right)$ is the Hard-Sigmoid activation function.
Finally, $f_{s}$ is upsampled and concatenated with $u_{1}$ to generate the resulting feature map $D_{1}$. The $D_{1}$ undergoes a standard convolution to restore the input resolution, followed by a 1 $\times$ 1 convolution for channel adjustment prior to final prediction.